OpenStudy (anonymous):

Medal + Fan: The slope-intercept form of the equation of a line that passes through point (–2, –13) is y = 5x – 3. What is the point-slope form of the equation for this line?

OpenStudy (haseeb96):

as we know that the given equation is y=5x-3 here m=5 is the slope while point is given that is (-2,-13) now put it in this formula and get your answer (y-y1) = m (x-x1) y-(-13) = 5(x-(-2)) y+13 = 5x +10 y=5x +10-13 y=5x -3 is your answer

OpenStudy (anonymous):

Hmm doesn't seem right, did you forget to subtract or add from both sides?

OpenStudy (anonymous):

These were the options: y – 13 = 5(x – 2) y + 13 = 5(x + 2) y – 2 = 5(x – 13) y + 2 = 5(x + 13)

OpenStudy (haseeb96):

B is the correct answer

OpenStudy (haseeb96):

i have written it in step # 3

OpenStudy (anonymous):

Oh could you explain why , I couldnt see it haha

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

I see what you did, thank you.

OpenStudy (haseeb96):

Do you have more questions ? @Roman532

OpenStudy (anonymous):

yes I do it would be really appreciated if you could help... much apprreciated.

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

I got 10 min remaining

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

Tickets to a basketball game can be ordered online for a set price per ticket plus a $5.50 service fee. The total cost in dollars for ordering 5 tickets is $108.00. Which linear function represents c, the total cost, when x tickets are ordered? c(x) = 5.50 + 20.50x c(x) = 5.50x + 20.50 c(x) = 5.50 + 21.60x c(x) = 5.50x + 21.60

OpenStudy (haseeb96):

can you post it with the new question?

OpenStudy (anonymous):

I got C?

OpenStudy (anonymous):

sure